the circuit below consists of a d.c supply of e.m.f of 45V and negligible internal resistance and three resistors. two of the resistors have resistances of 1000Ω and 750Ω as shown. the current drawn from the supply is 0.030A. Calculate the resistance of R



1
Expert's answer
2021-07-20T15:22:39-0400

Gives

Emf=45V

r=0 ohm

"R_1=1000 ohm" "R_2=750ohm"

"I=0.03A"


"R_{eq}=1000+750+R=(1750+R) ohm"

V=IR

Put value

"45=0.03\times(1750+R)"

"1750+R=1500"

R=(1750-1500 ) ohm=250"\Omega"